# AgriFM: A Multi-source Temporal Remote Sensing Foundation Model for Crop Mapping

### v2

@@ Abstract
Escalating climate change and global population growth demand precise crop mapping solutions to ensure food security. Accurate crop mapping fundamentally relies on modeling multi-scale spatiotemporal patterns, where spatial scales range from individual field textures to landscape-level context, and temporal scales capture both short-term phenological transitions and full growing-season dynamics, demanding robust joint spatiotemporal modeling capabilities. Existing approaches are inherently limited by their architectural designs that process spatial and temporal features separately, leading to significant information loss. Transformer-based remote sensing foundation models (RSFMs) offer promising potential for crop mapping due to their innate ability for unified spatiotemporal processing via self-attention mechanisms. However, current RSFMs remain suboptimal for crop mapping: they either employ fixed spatiotemporal windows that ignore the multi-scale nature of crop systems or completely disregard temporal information by focusing solely on spatial patterns. To bridge these gaps, we present AgriFM, a multi-source temporal remote sensing foundation model specifically designed for agricultural crop mapping. Our approach begins by establishing the necessity of simultaneous hierarchical spatiotemporal feature extraction, leading to the development of a modified Video Swin Transformer architecture where temporal down-sampling is synchronized with spatial down-sampling operations. This modified backbone enables efficient unified processing of long time-series satellite inputs while preserving critical multi-scale spatial patterns and phenological dynamics. AgriFM leverages temporally rich data streams from three satellite sources including MODIS, Landsat-8/9 and Sentinel-2, and is pre-trained on a global representative dataset comprising over 25 million image samples supervised by land cover products. The resulting framework incorporates a versatile decoder architecture that dynamically fuses these learned spatiotemporal representations, supporting diverse downstream tasks including cropland mapping, field-boundary delineation, early-season crop mapping, and specific crop mapping (e.g., winter wheat and paddy rice) with difference data sources. Comprehensive evaluations demonstrate AgriFM’s superior performance over conventional deep learning approaches and state-of-the-art general-purpose RSFMs across all downstream tasks. Notably, AgriFM demonstrates remarkable extensibility by supporting multi-source remote sensing data and variable temporal windows for diverse mapping tasks. Codes will be available at https://github.com/flyakon/AgriFM and https://glass.hku.hk.
@@
Foundation model, crop mapping, remote sensing, deep learning
@@ 1 Introduction
The dual challenges of rapid population growth and intensifying climate change impacts have elevated global food security to a pressing priority (Wheeler and Von Braun, 2013; Singh et al., 2023). Satellite remote sensing has become indispensable for addressing this challenge, providing effective tools for large-scale agricultural monitoring and precise crop mapping (Van Tricht et al., 2023; Fang et al., 2024; Chen et al., 2025b). Recent technological advances have enhanced both the spatial and temporal resolution of Earth observation data (Yuan et al., 2020), creating unprecedented opportunities for detailed crop mapping. However, it simultaneously demands more sophisticated analytical methods capable of extracting meaningful agricultural intelligence from complex spatiotemporal patterns.
@@
The growing demands for accurate crop mapping have spurred significant methodological evolution, beginning with phenology-based approaches. They utilize the distinct physical attributes of various crops, including their differing reflectance across spectral bands and their unique phenological traits at various growth phases. Examining and detecting the alterations throughout the crop growth cycle can efficiently differentiate various crops (Qiu et al., 2022; Dong et al., 2020; Qiu et al., 2017; Liang et al., 2024; Ashourloo et al., 2022). Qiu et al. (Qiu et al., 2022) introduced an innovative approach for generating yearly 500-m MODIS-derived national maps of various cropping systems in China, employing phenology-based mapping algorithms and pixel purity thresholds, resulting in an overall accuracy of 89%. Ashourloo et al. (Ashourloo et al., 2022) introduced a novel phenology-based approach utilizing Sentinel-2 time-series data for the efficient differentiation of wheat and barley across extensive regions, attaining an overall accuracy exceeding 76%. The issue with these methods is that they require setting different thresholds based on the mapping task and study area, which significantly affects their applicability. For complex scenarios and tasks, this may lead to erroneous or even failed mapping results.
@@
Machine learning methods (Yin et al., 2020; Yang et al., 2023; Xuan et al., 2023; Van Tricht et al., 2023) have, to a certain extent, resolved the issue of adaptability. They establish the relationship between satellite spectral data and crop types, utilizing algorithms such as Random Forest (Breiman, 2001) and Support Vector Machine (SVM) (Hearst et al., 1998). The effectiveness of machine learning largely depends on obtaining sufficient ground truth data, which is challenging in many scenarios. Van Tricht et al. (Van Tricht et al., 2023) introduced the WorldCereal, a global applicable, seasonally updated crop and irrigation mapping product, which leverages the Random Forest algorithm. He et al. (Yin et al., 2020) developed a novel method based on the Random Forest algorithm to map the extent and timing of abandoned cropland using the Landsat time series and tested this approach in 14 diverse global study regions. Despite the wide application in crop mapping tasks, machine learning methods struggle to simultaneously understand the spatial and temporal information from satellite observations. As a result, they often underperform in complex tasks and lack sufficient processing capacity for large-scale remote sensing data.
@@
Recently, transformer models have achieved breakthrough progress in computer vision (Zhang et al., 2024a), remote sensing image processing (Chen et al., 2024), and earth observation tasks (Li et al., 2024). Remote Sensing Foundation Models (RSFMs), built on the basis of transformers, are considered a superior paradigm for handling multi-source temporal remote sensing data (Zhou et al., 2023). Vision Transformer (ViT) (Dosovitskiy et al., 2020) and Swin Transformer (Liu et al., 2021, 2022) are two prominent transformer architectures utilized in the construction of foundation models. RSFMs typically use a large volume of remote sensing data for pre-training, and then a small amount of labeled data for fine-tuning to accomplish specific tasks (Tan et al., 2023; Zhu et al., 2024; Zhou et al., 2024; Lu et al., 2024; Zhang et al., 2024b). The key to pre-training is to help the network learn representations from a large amount of remote sensing data (Cong et al., 2022; Reed et al., 2023). Pre-training methods include masked imaging modeling (MIM) (Hondru et al., 2024) and contrastive learning (CL) (Chen et al., 2020; He et al., 2020; Hondru et al., 2024). In some cases, if a large amount of labeled remote sensing data can be obtained, supervised pre-training is also an effective pre-training method (Feng et al., 2021; Chen et al., 2025a).
Transformer-based RSFMs have demonstrated superior capabilities in extracting spatiotemporal features across various remote sensing tasks, significantly outperforming traditional machine learning methods and deep learning architectures. This suggests RSFMs hold substantial potential for crop mapping, as they can theoretically model complete crop growth cycles end-to-end while integrating multi-source remote sensing data to improve mapping accuracy. Several pioneering studies have begun exploring RSFMs for crop mapping tasks. Fang et al. (Fang et al., 2025) developed a rice mapping method using NASA-IBM’s Prithvi foundation model (Jakubik et al., 2023) with Harmonized Landsat and Sentinel-2 (HLS30) data (Claverie et al., 2018) and successfully generated rice distribution maps across Monsoon Asia. Similarly, Qi et al. (Qin et al., 2025) proposed a spatiotemporal masking strategy for pretraining a spatiotemporal collaborative learning network (STCLN) to extract informative representations for crop mapping.
@@
In contrast, Swin Transformer’s hierarchical feature extraction capability - resembling CNN’s multi-scale processing - appears better suited for crop mapping tasks. Like U-Net architectures, Swin Transformer can effectively fuse features across scales to enhance spatial precision (Cao et al., 2022). However, existing Swin-based RSFMs remain scarce and typically ignore temporal information, severely limiting their applicability for crop monitoring. While Video Swin Transformer (Liu et al., 2022) inherently supports joint spatiotemporal processing, few existing RSFMs have yet adopted it as the backbone architecture.
@@
To address these gaps, this paper first establishes the necessity of simultaneous hierarchical spatiotemporal feature extraction for crop mapping, leading to the development of a modified Video Swin Transformer (Liu et al., 2022) architecture where temporal down-sampling is synchronized with spatial down-sampling operations. This modified backbone enables efficient unified processing of long time-series satellite inputs while preserving critical multi-scale spatial patterns and phenological dynamics. We then develop AgriFM, a multi-source temporal remote sensing foundation model specifically designed for agricultural crop mapping.
@@
AgriFM leverages temporally rich data streams from MODIS, Landsat-8 / 9, and Sentinel-2 satellites, pre-trained on a global representative dataset with supervision from global land cover products. The pretraining dataset comprises over 25 million globally sampled images from MODIS (250m & 500m), Landsat-8/9 (30m), and Sentinel-2 (10m & 20m). For each satellite source, we randomly selected 16 chronological frames per year to capture complete seasonal variations. Unlike RSFMs relying on MIM or CL for pretraining, we incorporate land cover fraction from GLC_FCS30D data (Zhang et al., 2024c) for supervised pretraining. Specifically, we extract image-level land cover fractions within each sample area as regression targets, optimizing with L1 loss. This approach follows established practices demonstrating land cover priors’ effectiveness (Li et al., 2022; Wang et al., 2024). To address inherent label noise in reference products, we implemented an auxiliary teacher network with exponential moving average updates, effectively guiding the model toward robust feature learning while filtering out unreliable supervision signals.
@@
AgriFM incorporates a versatile decoder to dynamically fuse extracted multi-scale spatiotemporal representations, supporting diverse crop mapping tasks, including cropland mapping, field-boundary delineation, early-season crop mapping and specific crop mapping (e.g., winter wheat and paddy rice). For systematic evaluation, we compare AgriFM with three representative approaches: (1) ViT-based temporal RSFMs (Prithvi (Jakubik et al., 2023), SatMAE (Cong et al., 2022)), (2) Swin-based models without temporal processing (PIS (An et al., 2024)), and (3) conventional deep learning methods (CNNs / LSTMs). Experimental results demonstrate AgriFM’s consistent superiority, particularly in two crucial aspects: preserving fine-grained spatial details essential for field-level mapping, and effectively modeling long-term temporal patterns critical for crop type analysis - capabilities where existing methods show notable limitations. AgriFM achieves unified processing of heterogeneous data sources, variable temporal sequences, and diverse mapping tasks within a single network architecture. This comprehensive method overcomes the traditional need for task-specific model adaptations while delivering robust performance across various crop mapping tasks.
@@ 2 Methodology and Data
In this section, we will delve into the implementation specifics surrounding the AgriFM. The entire flowchart is graphically represented in Figure 1 and can be systematically divided into two primary phases. The first phase involves the creation of a large-scale pre-training dataset and the extraction of land cover fractions to serve as pre-training supervision. The second phase pertains to the development of our multi-source temporal foundation model, along with details of the pre-training process. Based on the foundation mode AgriFM, a unified crop mapping framework is achieved through the construction of a versatile decoder and subsequent fine-tuning with labels.
Figure 1: Flowchart outlining the AgriFM: a) The initial phase involves the extraction of pre-training supervision from geographical priors (land cover products) and the assembly of an extensive pre-training dataset, b) The subsequent phase entails the pre-training of the multi-source temporal foundation model and construction of unified crop mapping framework.
@@ 2.1.1 Multi-source Satellite Data for Pre-training
The MODIS data from the Terra satellite include the surface reflectance products MOD09A1 and MOD09Q1 with the temporal resolution of 8 days and the spatial resolutions of 500 m and 250 m, respectively. The MOD09Q1 product comprises the first two bands at 250m, and we extract bands 3-7 at 500m from the MOD09A1 product and then re-sample them to 250 meters. The MODIS data selected for pretraining are from 2020 and 2021.
@@ 2.1.2 Land Cover Product for Pre-training Supervision
During the pre-training phase, we segment the original satellite data into 224 x 224 sized images. For each image, we determine the land cover distribution within the region it covers, and calculate the fractions accordingly (Li et al., 2021). These land cover fractions will then serve as pre-training supervision during the pre-training phase. For any input remote sensing image $\mathbf{I}$, we first extract its geographic coordinate boundaries defined as $\Omega=[x_{min},x_{max}]\times[y_{min},y_{max}]$, where $x_{min}$, $x_{max}$, $y_{min}$, and $y_{max}$ represent the minimum longitude, maximum longitude, minimum latitude, and maximum latitude of the image footprint, respectively. From the land cover product $\mathcal{L}$, we retrieve the corresponding classification map $\mathbf{L}$ within $\Omega$:
@@
where $\mathbf{L}_{i,j}\in{1,2,…,K}$ denotes the land cover class at pixel $(i,j)$, with $K$ being the total number of land covers. In our implementation, we focus on the eight most frequent land covers in the dataset: (1) cropland, (2) forest, (3) shrubland, (4) grassland, (5) wetland, (6) water, (7) bare land, and (8) urban areas. Pixels belonging to other classes are aggregated as background.
The land cover fraction $f_{k}$ for the k-th land cover class within $\Omega$ is computed as:
$p_{k}=\frac{1}{N}\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbb{I}(L_{ij}=k).$ (2)
@@
where $p_{0}$ represents the background land cover and $p_{1}$ to $p_{8}$ represent the fraction of each land cover.

### v3

# AgriFM: A Multi-source Temporal Remote Sensing Foundation Model for Agriculture Mapping

@@ Abstract
Climate change and population growth intensify the demand for precise agriculture mapping to enhance food security. Such mapping tasks require robust modeling of multi-scale spatiotemporal patterns from fine field textures to landscape context, and from short-term phenology to full growing-season dynamics. Existing methods often process spatial and temporal features separately, limiting their ability to capture essential agricultural dynamics. While transformer-based remote sensing foundation models (RSFMs) offer unified spatiotemporal modeling ability, most of them remain suboptimal: they either use fixed windows that ignore multi-scale crop characteristics or neglect temporal information entirely. To address these gaps, we propose AgriFM, a multi-source, multi-temporal foundation model for agriculture mapping. AgriFM introduces a synchronized spatiotemporal downsampling strategy within a Video Swin Transformer backbone, enabling efficient handling of long and variable-length satellite time series while preserving multi-scale spatial and phenological information. It is pre-trained on a globally representative dataset comprising over 25 million samples from MODIS, Landsat-8/9, and Sentinel-2 with land cover fractions as pre-training supervision. AgriFM further integrates a versatile decoder specifically designed to dynamically fuse multi-source features from different stages of backbone and accommodate varying temporal lengths, thereby supporting consistent and scalable agriculture mapping across diverse satellite sources and task requirements. It supports diverse tasks including agricultural land mapping, field boundary delineation, agricultural land use / land cover mapping, and specific crop mapping (e.g., winter wheat and paddy rice) with difference data sources. Comprehensive evaluations show that AgriFM consistently outperforms existing deep learning models and general-purpose RSFMs across multiple agriculture mapping tasks. Codes and models are available at https://github.com/flyakon/AgriFM and https://glass.hku.hk.
@@
Foundation model, agriculture mapping, remote sensing, deep learning
@@ 1 Introduction
The dual challenges of rapid population growth and intensifying climate change impacts have elevated global food security to a pressing priority (Wheeler and Von Braun, 2013; Singh et al., 2023). Satellite remote sensing has become indispensable for addressing this challenge, providing effective tools for large-scale agricultural monitoring and precise crop mapping (Van Tricht et al., 2023; Fang et al., 2024; Chen et al., 2025b). Recent technological advances have enhanced both the spatial and temporal resolution of Earth observation data (Yuan et al., 2020; Ma et al., 2025; Liang et al., 2021; Jiang et al., 2023), creating unprecedented opportunities for detailed agriculture mapping. However, it simultaneously demands more sophisticated analytical methods capable of extracting meaningful agricultural intelligence from complex spatiotemporal patterns.
@@
The growing demands for accurate agriculture mapping (e.g. agricultural land mapping, agricultural land use / land cover mapping, and crop mapping) have spurred significant methodological evolution, beginning with phenology-based approaches. They utilize the distinct physical attributes of various crops, including their differing reflectance across spectral bands and their unique phenological traits at various growth phases. Examining and detecting the alterations throughout the crop growth cycle can efficiently differentiate various crops (Qiu et al., 2022; Dong et al., 2020; Qiu et al., 2017; Liang et al., 2024; Ashourloo et al., 2022). Qiu et al. (Qiu et al., 2022) introduced an innovative approach for generating yearly 500-m MODIS-derived national maps of various cropping systems in China, employing phenology-based mapping algorithms and pixel purity thresholds, resulting in an overall accuracy of 89%. Ashourloo et al. (Ashourloo et al., 2022) introduced a novel phenology-based approach utilizing Sentinel-2 time-series data for the efficient differentiation of wheat and barley across extensive regions, attaining an overall accuracy exceeding 76%. The issue with these methods is that they require setting different thresholds based on the mapping task and study area, which significantly affects their applicability. For complex scenarios and tasks, this may lead to erroneous or even failed mapping results.
@@
Machine learning methods (Yin et al., 2020; Yang et al., 2023; Xuan et al., 2023; Van Tricht et al., 2023) have, to a certain extent, resolved the issue of adaptability. They establish the relationship between satellite spectral data and agriculture land use / land covers, utilizing algorithms such as Random Forest (Breiman, 2001) and Support Vector Machine (SVM) (Hearst et al., 1998). The effectiveness of machine learning largely depends on obtaining sufficient ground truth data, which is challenging in many scenarios. Van Tricht et al. (Van Tricht et al., 2023) introduced the WorldCereal, a global applicable, seasonally updated crop and irrigation mapping product, which leverages the Random Forest algorithm. Yin et al. (Yin et al., 2020) developed a novel method based on the Random Forest algorithm to map the extent and timing of abandoned cropland using the Landsat time series and tested this approach in 14 diverse global study regions. Despite the wide application in agriculture mapping tasks, machine learning methods struggle to simultaneously understand the spatial and temporal information from satellite observations. As a result, they often underperform in complex tasks and lack sufficient processing capacity for large-scale remote sensing data.
@@
Recently, transformer models have achieved breakthrough progress in computer vision (Zhang et al., 2024a), remote sensing image processing (Chen et al., 2024), and earth observation tasks (Li et al., 2024). Remote Sensing Foundation Models (RSFMs), built on the basis of transformers, are considered a superior paradigm for handling multi-source temporal remote sensing data (Zhou et al., 2024a). Vision Transformer (ViT) (Dosovitskiy et al., 2021) and Swin Transformer (Liu et al., 2021, 2022) are two prominent transformer architectures utilized in the construction of foundation models. RSFMs typically use a large volume of remote sensing data for pre-training, and then a small amount of labeled data for fine-tuning to accomplish specific tasks (Tan et al., 2023; Zhu et al., 2024; Zhou et al., 2024b; Lu et al., 2024; Zhang et al., 2024b). The key to pre-training is to help the network learn representations from a large amount of remote sensing data (Cong et al., 2022; Reed et al., 2023). Pre-training methods include masked imaging modeling (MIM) (Hondru et al., 2025) and contrastive learning (CL) (Chen et al., 2020; He et al., 2020; Hondru et al., 2025). In some cases, if a large amount of labeled remote sensing data can be obtained, supervised pre-training is also an effective pre-training method (Feng et al., 2021; Chen et al., 2025a).
Transformer-based RSFMs have shown strong potential in extracting spatiotemporal features from remote sensing data. This ability to model temporal sequences end-to-end and integrate multi-source data makes them particularly promising for agricultural applications, where capturing complete crop growth cycles is critical for accurate mapping(Tseng et al., 2025; Sumbul et al., 2025). Several pioneering studies have begun exploring RSFMs for crop mapping tasks. Fang et al. (Fang et al., 2025) developed a rice mapping method using NASA-IBM’s Prithvi foundation model (Jakubik et al., 2023) with Harmonized Landsat and Sentinel-2 (HLS30) data (Claverie et al., 2018) and successfully generated rice distribution maps across Monsoon Asia. Similarly, Qin et al. (Qin et al., 2025) proposed a spatiotemporal masking strategy for pretraining a spatiotemporal collaborative learning network (STCLN) to extract informative representations for crop mapping.
@@
In contrast, Swin Transformer’s hierarchical feature extraction capability, resembling CNN’s multi-scale processing, appears better suited for agriculture mapping tasks. Like U-Net architectures, Swin Transformer can effectively fuse features across scales to enhance spatial precision (Cao et al., 2022). However, existing Swin-based RSFMs remain scarce and typically ignore temporal information, severely limiting their applicability for crop monitoring. While Video Swin Transformer (Liu et al., 2022) inherently supports joint spatiotemporal processing, few existing RSFMs have yet adopted it as the backbone architecture.
@@
To address these gaps, this paper first establishes the necessity of simultaneous hierarchical spatiotemporal feature extraction for agriculture mapping, leading to the development of a modified Video Swin Transformer (Liu et al., 2022) architecture where temporal down-sampling is synchronized with spatial down-sampling operations. This modified backbone enables efficient unified processing of long and variable-length satellite time series while preserving critical multi-scale spatial patterns and phenological dynamics. We then develop AgriFM, a multi-source temporal remote sensing foundation model specifically designed for agriculture mapping.
AgriFM leverages temporally rich data streams from MODIS, Landsat-8 / 9, and Sentinel-2 satellites, pre-trained on a global representative dataset with supervision from global land cover products. The pretraining dataset comprises over 25 million globally sampled images from MODIS (250m & 500m), Landsat-8/9 (30m), and Sentinel-2 (10m & 20m). Leveraging the synchronized spatio-temporal downsampling strategy, we dynamically sample 3-32 frames from each satellite source during every pre-training iteration. By adaptively adjusting the downsampling ratios across temporal scales, feature maps of consistent dimensions can be obtained while accommodating variable sequence lengths. The foundation model thus learns to extract robust representations from diverse temporal contexts, effectively capturing multi-scale phenological patterns across different agricultural monitoring scenarios.
@@
Unlike other RSFMs relying on MIM or CL for pretraining, we incorporate land cover fraction from GLC_FCS30D data (Zhang et al., 2024c) for supervised pretraining. Specifically, we extract image-level land cover fractions within each sample area as regression targets, optimizing with L1 loss. This approach follows established practices demonstrating land cover priors’ effectiveness (Li et al., 2022; Wang et al., 2024). To mitigate the impact of supervision noises caused by inaccuracies in reference products, we implement an auxiliary teacher network with exponential moving average updates, effectively guiding the model toward robust feature learning while filtering out unreliable supervision signals.
@@
AgriFM incorporates a versatile decoder to dynamically fuse extracted multi-scale spatiotemporal representations, supporting diverse agriculture mapping tasks, including agricultural land mapping, field boundary delineation, agricultural land use/ land cover mapping and specific crop mapping (e.g., winter wheat and paddy rice). For systematic evaluation, we compare AgriFM with three representative approaches: (1) ViT-based RSFMs (Prithvi (Jakubik et al., 2023), SatMAE (Cong et al., 2022), Galileo (Tseng et al., 2025), SMARTIES (Sumbul et al., 2025)), (2) Swin-based models without temporal processing (PIS (An et al., 2024), GFM (Mendieta et al., 2023)), and (3) conventional deep learning methods (CNNs / LSTMs). Experimental results demonstrate AgriFM’s consistent superiority, particularly in two crucial aspects: preserving fine-grained spatial details essential for field-level mapping, and effectively modeling long-term temporal patterns critical for crop type analysis, where existing methods show notable limitations. AgriFM achieves unified processing of heterogeneous data sources, variable temporal sequences, and diverse mapping tasks within a single network architecture. This comprehensive method overcomes the traditional need for task-specific model adaptations while delivering robust performance across various tasks.
@@ 2 Methodology and Data
In this section, we will delve into the implementation specifics surrounding the AgriFM. The entire flowchart is graphically represented in Figure 1 and can be systematically divided into two primary phases. The first phase involves the creation of a large-scale pre-training dataset and the extraction of land cover fractions to serve as pre-training supervision. The second phase pertains to the development of our multi-source temporal foundation model, along with details of the pre-training process. Based on the foundation mode AgriFM, a unified mapping framework is achieved through the construction of a versatile decoder and subsequent fine-tuning with labels.
Figure 1: Flowchart outlining the AgriFM: a) The initial phase involves the extraction of pre-training supervision from geographical priors (land cover products) and the assembly of an extensive pre-training dataset, b) The subsequent phase entails the pre-training of the multi-source temporal foundation model and construction of unified mapping framework.
@@ 2.1.1 Multi-source Satellite Data for Pre-training
The MODIS data from the Terra satellite include the surface reflectance products MOD09A1 and MOD09Q1 with the temporal resolution of 8 days and the spatial resolutions of 500 m and 250 m, respectively. The MOD09Q1 product comprises the first two bands at 250m, and we extract bands 3-7 at 500m from the MOD09A1 product and then re-sample them to 250 meters. The MODIS data selected for pretraining are from 2020 to 2022.
@@ 2.1.2 Land Cover Product for Pre-training Supervision
During the pre-training phase, we segment the original satellite data into 224 x 224 sized images. For each image, we determine the land cover distribution within the region it covers, and calculate the fractions accordingly (Li et al., 2021). These land cover fractions will then serve as pre-training supervision during the pre-training phase. For any input remote sensing image $\mathbf{I}$, we first define its geographic footprint. We denote the coordinates of the upper left corner as $lon_{min},lat_{max}]$ and the lower right corner as $lon_{max},lat_{min}]$. The spatial extent $\Omega$ is then defined as the rectangular region spanning from $lon_{min},lat_{max}]$ to $lon_{max},lat_{min}]$. From the land cover product $\mathcal{L}$, we retrieve the corresponding classification map $\mathbf{L}$ within $\Omega$:
@@
where $\mathbf{L}$ denotes the land cover map and $\mathbf{L}_{i,j}\in{1,2,…,K}$ denotes the land cover class at pixel $(i,j)$, with $K$ being the total number of land covers. From this step, we extract all pixels in region $\Omega$ with upper left coordinates $lon_{min},lat_{max}]$ and lower right coordinates $lon_{max},lat_{min}]$ In our implementation, we focus on the eight most frequent land covers in the dataset: (1) cropland, (2) forest, (3) shrubland, (4) grassland, (5) wetland, (6) water, (7) bare land, and (8) urban areas. Pixels belonging to other classes are aggregated as background.
The land cover fraction for the k-th land cover class within $\Omega$ is denoted as $p_{k}$ and is computed as:
$p_{k}=\frac{1}{N}\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbb{I}(\mathbf{L}_{i,j}=k).$ (2)
@@
where $p_{1}$ to $p_{8}$ represent the fractional coverage of their respective land cover classes, while $p_{0}$ denotes the collective fraction of all other categories not explicitly listed, serving as the background or no-data component.